OpenStudy (anonymous):

Solve 2/3 + n + 6 = 3/4

OpenStudy (anonymous):

\[\frac{2}{3} + n + 6 = \frac{3}{4}\]\[\frac{2}{3} + n + \frac{18}{3} = \frac{3}{4}\]\[\frac{20}{3} + n = \frac{3}{4}\]\[n = \frac{3}{4}-\frac{20}{3}\]\[n = \frac{9}{12}-\frac{80}{12}\]\[n =-\frac{71}{12}\]

OpenStudy (anonymous):

thanks for the great workout, now I understand :-)